# Junyu Cao and Mohsen Bayati

### v1

@@ 5 Practical Implementations
In the preceding sections, we established a framework that assumes humans can compare any two parametric models. However, such a comparison could be possibly challenging when humans only have access to the models themselves. In practice, it is often more feasible to compare models based on their performance with sampled data, allowing humans to discern which one aligns better with the true underlying model. To address this specific challenge, this section delves into the practical aspects of comparing two models using selected samples.
@@ 5.1 Active Learning: How to Select Samples for Comparison
Active learning focuses on the strategic selection of samples to be labeled or compared in our context, with the aim of constructing prediction models in a resource-efficient manner. The key idea is to assess the significance of a sample to determine the value of acquiring its labels. Algorithm 2 actively selects a pair of two models to ask for comparison. In practical scenarios, it often involves soliciting human judgment to compare the responses predicted by two different models for a given sample data, rather than directly comparing their model parameters.
We start with an example. The clinician needs to evaluate the Atherosclerotic Cardiovascular Disease (ASCVD) risk of patients $f_{\boldsymbol{\theta}}$, which is a linear function of contextual information, denoted as $x$, such as demographics (age, gender, racial/ethnic group, and education level), baseline conditions (body mass index, history of ASCVD events, and smoking status), clinical variables (glycosylated hemoglobin (A1c) and systolic blood pressure (sbp)), anti-hypertensive agenets, anti-hyperglycemic agents.
Given sample of patient context $x\in\mathcal{X}$, the clinician evaluates the ASCVD risk according to their expertise. The distance between model $f_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}^{*}}$ evaluated on sample $x$ can be defined as the absolute value of the difference:
@@
Then the utility of $f_{\boldsymbol{\theta}}$ on sample $x$ is $u_{x}(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*}|x)=-|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|$. When presented with risk levels predicted by $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the clinician would typically opt for the one that is closer to their assessment.
Let us first consider a simple scenario. When $\mathcal{X}$ contains adequate variety of samples such that for any base vector $e_{i}$, there exists sample $x$ such that $\varphi(x)=e_{i}$, then by comparing $f_{\boldsymbol{\theta}_{1}}$ with $f_{\boldsymbol{\theta}_{2}}$ on such sample $x$, we have
@@
It implies that the comparison of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ when testing on $x$ would be solely on the $i^{th}$ dimension. Thus, the alignment process for the parameter in different dimensions can be decoupled, and SFT+HC can be directly applied where the Stage 2 alignment is proceeded along each dimension.
However, in practice, the dataset $\mathcal{X}$ may be limited, potentially leading to situations where there are no samples $x\in\mathcal{X}$ that satisfy the condition $\varphi(x)=e_{i}$. In such cases, it becomes impractical to refine the model parameter along each dimension separately. The question then arises: how can we efficiently select samples to refine the model parameter? To address this challenge, we propose a procedure for constructing a new basis using the available samples in $\mathcal{X}$. This procedure aims to optimize the alignment process by leveraging the existing data to guide the selection of samples for parameter enhancement.
5.2 New Basis Construction through Gram Schmidt Process
To facilitate the independent alignment of each coordinate, we aim to identify a new set of orthogonal basis, which can be constructed using $\Psi:=\{\varphi(x):x\in\mathcal{X}\}$. Let $\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}$ represent these new basis, and our objective is to learn $\boldsymbol{\theta}^{*}=\sum_{i=1}^{s}\omega_{i}^{*}\boldsymbol{\alpha}_{i}$, where $\omega_{i}^{*}$ are the corresponding coefficients. We describe the process of constructing the orthogonal basis $\{\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}\}$ from the sample space.
First, we pick a set of samples that spans the space of $\Psi$, denoted as $\varphi(x_{1}),\cdots,\varphi(x_{s})$. For simplicity, we define $\boldsymbol{z}_{i}=\varphi(x_{i})$. Then we use the Gram Schmidt process to construct a set of orthogonal basis based on $\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s}$:
@@ 5.2 New Basis Construction through Gram Schmidt Process
We sequentially align the coordinates $\omega_{1},\cdots,\omega_{s}$. When evaluating on sample $x_{1}$, the true value is represented by $\varphi(x_{1})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{1}^{\top}\boldsymbol{\theta}^{*}=\omega_{1}^{*}\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1}$. Utilizing a probabilistic bisection process in 2, we actively select two answers for each query. Through this iterative refinement, the final value terminates at $y(x_{1})$, which is guaranteed to be within $\varepsilon$-accuracy of the true parameter $\omega_{1}^{*}$ (with high probability). In the end, the first coordinate is refined to $\hat{\omega}_{1}=y(x_{1})/(\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1})$.
@@
where $w_{1}$ is refined already. Once more, employing the probabilistic bisection process, the value of $z_{2}^{\top}\boldsymbol{\theta}^{*}$ is iteratively refined to $y(x_{2})$. Upon termination of the probabilistic bisection process, we obtain the aligned value of the second coordinate as
@@
Subsequently, at the $k^{th}$ iteration, we test on the sample $\boldsymbol{z}_{k}=\boldsymbol{\alpha}_{k}+\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$ where the true value equals
@@
Once the value of $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}$ is refined to $y(x_{k})$ through 2, we are able to obtain the aligned value of the $k^{th}$ coordinate as
@@
Through asking human to evaluate $x_{1},\cdots,x_{s}$ and refining $\omega_{1},\cdots,\omega_{s}$ in sequence, the parameter $\boldsymbol{\theta}:=\sum_{i=1}^{s}\hat{\omega}_{i}\boldsymbol{\alpha}_{i}$ is refined within a certain accuracy level. We include the pseudo code in Algorithm 4.
At the end, it is essential to emphasize that the set of orthogonal basis is not unique, as there may exist multiple sets. Consequently, we can conduct multiple rounds of alignment, with the option to switch basis from one round to another.
@@
6:     Query sample $x_{k}$ through 2 and get output $y(x_{k})$;
@@ 6 Numerical Experiment
6.1 Testing Sample Complexity
In the first experiment, we compare the performance of SFT+HC and pure SFT under the same sample size.
Experiment setup. Consider a high-dimensional linear regression problem where $d=100$ and $s=10$. The first ten coordinates of the true parameter are independently drawn from the uniform distribution and the rest coordinates are zero. The observational noise is drawn from Gaussian distribution with mean zero and variance $\sigma=\{1,2,5\}$; we choose $\kappa\in[0,0.95]$. The selection precision (along each dimension) is set as $\varepsilon=0.1$ and the human expert level in the choice model is $\gamma=1$. We conducted 50 repetitions for each experiment and calculated the average across all results.
In the first stage the two-stage framework, we use lasso and select the sample size large enough such that the estimator for the non-zero coefficient can be deviated from zero. That is, we guarantee that important features are selected with high probability. Once the algorithm SFT+HC stops, we use the same sample size to train lasso and compute the estimation error. We compare the estimation error of two algorithms under the same sample size.
@@ 6.1 Testing Sample Complexity
Figure 3 shows the ratio of the estimation error of two algorithms (Err(SFT+HC)/Err(SFT)) with different values of $\kappa$ when $\sigma=1$. For smaller value of $\kappa$, human’s selection accuracy is higher when $\theta$ approaches $\theta^{*}$. In particular, when $\kappa=0$, the selection accuracy is bounded away from 50%. Therefore, the task is harder when $\kappa$ increases. It is consistent with the increasing trend of the estimation error when $\kappa$ grows from 0 to 1. In addition, a larger value of $\kappa$ leads to a larger sample size, so the estimation error of lasso decreases. The intersection point is around $\kappa=0.3$. When $\kappa\leq 0.3$, SFT+HC performs better than pure SFT. In particular, when $\kappa=0$, the error of two-stage framework is roughly 50% of that in SFT.
Figure 3: Error ratio of the two-stage framework/pure SFT when $\sigma=1$.
The results for noise levels with $\sigma=2$ and $\sigma=5$ are illustrated in Figures 4. Notably, the trends of the error curves exhibit similarity. The threshold, defined as the intersection point of these curves, is approximately 0.55 when $\sigma=2$. However, when $\sigma=5$, the two curves do not intersect, suggesting that SFT+HC consistently outperforms pure SFT.
A comparison across the three figures reveals that the superiority of SFT+HC becomes more pronounced in the presence of higher observational noise. In conclusion, SFT+HC is more effective under conditions of increased noise and reduced value of $\kappa$.
(a) Error ratio when $\sigma=2$.
(b) Error ratio when $\sigma=5$.
Figure 4: Prediction errors with different noise levels.
6.1.2 The variation of expert level $\gamma$.
We change the value of human expert level $\gamma$ to see how the human comparison accuracy would influence the estimation error. The smaller value of $\gamma$ indicates higher accuracy in the comparison task. In particular, when $\gamma$ approaches 0, human can always select the accurate answer. Figure 5 shows the performance of two algorithms with $\gamma=1.1$ and $\gamma=0.8$, respectively. Comparing two figures, when $\gamma=1.1$, SFT+HC performs better than pure SFT when $\kappa$ is lower than approximately 0.45; when $\gamma=0.8$, SFT+HC performs better than pure SFT when $\kappa$ is lower than approximately 0.65. Therefore, it is consistent with our theory that SFT+HC improves its performance when human has a higher expert level in the selection (i.e., the selection accuracy is higher).
(a) Error ratio when $\gamma=1.1$.
(b) Error ratio when $\gamma=0.8$.
Figure 5: Prediction errors with different expert levels.
@@
The relative performance of two algorithms also depends on the dimension of the problem. For the experiment setup, we set $\gamma=1$ and $\sigma=2$. Figure 6 illustrates the algorithmic performance with important feature dimension $s=20$ and $s=50$, respectively. The figure reveals that the intersection point shifts to a smaller value as the important feature dimension increases. This observation aligns with our theoretical condition (4.2), indicating that as the important feature dimension grows, the advantage of SFT+HC diminishes.
(a) Error ratio when $s=20$.
(b) Error ratio when $s=50$.
Figure 6: Prediction errors with different important feature dimensions.
6.2 Testing LNCA Ratio
The LNCA ratio, $\sigma^{2}/s^{\kappa}$, determines the benefits of SFT+HC over pure SFT. The LNCA ratio varies between different tasks. We are interested in estimating this value for a specific task: to determine the percentage of total area enclosed by a given (red) polygon, as shown in Figure 7, for example.
@@ 6.2 Testing LNCA Ratio
We divided participants into two distinct groups. The first group was presented with the polygon figure and asked to provide their estimated percentage of the total enclosed area. The second group was shown the same figure but presented with two options for the area percentage. Their task was to select the option they believed was closer to the actual percentage of the enclosed area. For instance, in Figure 7(b), the true percentage is 50%. Human is provided with two options, 71% and 55%, so the correct choice is 55%.
@@
This dual-group approach aimed to explore the efficacy of human estimation when directly asked to estimate the area versus when presented with comparative options. In the latter case, we varied the discrepancy between the two choices to assess how the ease or difficulty of comparison affected the accuracy of their selections. A larger disparity between the choices facilitated easier decision-making, while closer options posed a more challenging comparative task. Within Figure 7, three testing samples are depicted. Notably, sample (a) presents the most challenging comparison among these three instances due to the minimal disparity observed between them.
Fifty different tasks are launched in Amazon Mechanical Turk [^4] https://www.mturk.com/, as shown in Figure 8, and we collected 1300 data points in total. We use the following ways to estimate $\kappa$ and $\sigma$.
(a) 21% vs. 5%. True: 10%.
(b) 71% vs. 55%. True: 50%.
(c) 29% vs. 13%. True: 26%.
Figure 7: Experiment design.
(a) The first group: Select a better answer.
(b) The second group: Estimate the percentage.
Figure 8: Comparison vs. Estimation.
@@
Estimation process for $\kappa$. Given figures with true parameter values $\theta^{*}$ representing the percentage of total area enclosed by a polygon, we aim to estimate the accuracy convergence rate $\kappa$ between two options using a human choice model. We collect multiple samples for each figure and then use these estimates to derive the relationship between the difference in utilities and the accuracy of human choices.
@@
More specifically, among two choices, assume the smaller one is $c^{-}$ and the larger one is $c^{+}$, where the middle point is $\theta=(c^{-}+c^{+})/2$. According to the choice model, human makes the correct choice with accuracy $p(\theta)=\frac{1}{1+\exp(-|u(c^{-})-u(c^{+})|/\gamma)}$. Note that we try to estimate $\kappa$ from Assumption 3.4.1 that
@@
That is, the accuracy is $p(\theta)=\frac{1}{1+\exp(-\lambda_{\Delta}/\gamma\cdot|\theta-\theta^{*}|^{\kappa})}$. Let $\tilde{\lambda}=-\lambda_{\Delta}/\gamma$. For each data point with tested point $\theta_{i}$ and true parameter $\theta_{i}^{*}$, we define $y_{i}=1$ (the correct choice) with probability $p_{i}(\theta)$ and $y_{i}=0$ (the wrong choice) with probability $1-p_{i}(\theta)$. By maximizing the log-likelihood, we solve the optimization problem
$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})$
@@
$\max\quad\sum_{i=1}^{K}-y_{i}\log(1+\exp(-\tilde{\lambda}|\theta_{i}-\theta^{*}|^{\kappa}))-(1-y_{i})\log(1+\exp(\tilde{\lambda}|\theta_{i}-\theta_{i}^{*}|^{\kappa}))\,.$
From the data collected from Amazon Mechanical Turk, the estimated $\kappa$ is 0.095.
@@
Estimation process for $\sigma$. The other group is asked to directly estimate the percentage of total area. For the figure with true parameter values $\theta_{i}^{*}$, human provides responses by $y_{i}=\theta_{i}^{*}+\epsilon_{i}$. We estimate the variance of the noise by $\hat{\sigma}^{2}=\sum_{i=1}^{N}\frac{(y_{i}-\theta_{i}^{*})^{2}}{N-1}=29.04$.
In the context of a one-dimensional problem, as evidenced by numerical experiments detailed in the preceding section, SFT+FTPB demonstrates a substantial performance advantage over SFT when $\kappa\approx 0.1$ and $\sigma\approx 30$.
7 Concluding Remarks

### v2

@@ 6 Practical Implementations
In the preceding sections, we established a framework that assumes the opportunity for humans to compare any two parametric models. However, such a comparison could be challenging when humans have access only to the models themselves. In practice, comparing models based on their performance with sampled data is often more useful because this engagement allows humans to discern which model aligns better with the true underlying model. To address this specific challenge, in this section, we present the practical aspects of comparing two models using selected samples.
@@ 6.1 Active Learning: How to Select Samples for Comparison
Active learning focuses on the strategic selection of samples to be labeled or compared in our context, with the aim of constructing prediction models in a resource-efficient manner. The key idea is to assess the significance of a sample to determine the value of acquiring its labels. Algorithm 2 actively selects a pair of models for comparison. In practical scenarios, this choice often involves soliciting human judgment to compare the responses predicted by two different models for the given sample data, rather than directly comparing their model parameters.
We start with an example. The clinician needs to evaluate the atherosclerotic cardiovascular disease (ASCVD) risk of patients $f_{\boldsymbol{\theta}}$, which is a linear function of contextual information, denoted as $x$. Such information comprises demographics (age, gender, racial/ethnic group, and education level); baseline conditions (body mass index, history of ASCVD events, and smoking status); clinical variables (glycosylated hemoglobin (A1c) and systolic blood pressure (sbp)); anti-hypertensive agents; and anti-hyperglycemic agents.
Given a sample of patients’ contexts $x\in\mathcal{X}$, clinicians evaluate the ASCVD risk according to their expertise. The distance between models $f_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}^{*}}$, evaluated on sample $x$, can be defined as the absolute value of the difference:
@@
Then, the utility of $f_{\boldsymbol{\theta}}$ on sample $x$ is $u_{x}(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*}|x)=-|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|$. When presented with risk levels predicted by $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, clinicians would typically opt for the model that produces results closest to their assessment.
Let us first consider a simple scenario. When $\mathcal{X}$ contains an adequate variety of samples, such that for any base vector $e_{i}$, there exists sample $x$ such that $\varphi(x)=e_{i}$, then by comparing $f_{\boldsymbol{\theta}_{1}}$ with $f_{\boldsymbol{\theta}_{2}}$ on this sample $x$, we have
@@
It implies that the comparison of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$, when testing on $x$, would focus only on the $i^{th}$ dimension. Thus, the alignment process for the parameter in different dimensions can be decoupled, and SL+LHF can be directly applied, where the Stage 2 alignment proceeds along each dimension.
However, in practice, the dataset $\mathcal{X}$ may be limited, potentially leading to situations where there are no samples $x\in\mathcal{X}$ that satisfy the condition $\varphi(x)=e_{i}$. In such cases, refining the model parameter along each dimension separatelybecomes impractical. The question then arises: How can we efficiently select samples to refine the model parameter? To address this challenge, we propose a procedure for constructing a new basis of comparison using the available samples in $\mathcal{X}$. This procedure aims to optimize the alignment process by leveraging the existing data to guide the selection of samples for parameter enhancement.
6.2 New Basis Construction through the Gram Schmidt Process
To facilitate the independent alignment of each coordinate, we aim to identify a new set of orthogonal basis, which can be constructed using $\Psi:=\{\varphi(x):x\in\mathcal{X}\}$. Let $\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}$ represent these new bases, and our objective is to learn $\boldsymbol{\theta}^{*}=\sum_{i=1}^{s}\omega_{i}^{*}\boldsymbol{\alpha}_{i}$, where $\omega_{i}^{*}$ are the corresponding coefficients. We describe the process of constructing the orthogonal basis $\{\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}\}$ from the sample space.
First, we pick a set of samples that spans the space of $\Psi$, denoted as $\varphi(x_{1}),\cdots,\varphi(x_{s})$. For simplicity, we define $\boldsymbol{z}_{i}=\varphi(x_{i})$. Then, we use the Gram Schmidt process to construct a set of orthogonal bases, based on $\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s}$:
@@ 6.2 New Basis Construction through the Gram Schmidt Process
We sequentially align the coordinates as $\omega_{1},\cdots,\omega_{s}$. When evaluating a model on sample $x_{1}$, the true value is represented by $\varphi(x_{1})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{1}^{\top}\boldsymbol{\theta}^{*}=\omega_{1}^{*}\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1}$. Using a probabilistic bisection process in Algorithm 2, we actively select two answers for each query. Through this iterative refinement, the final value terminates at $y(x_{1})$, which is guaranteed to be within $\varepsilon$-accuracy of the true parameter $\omega_{1}^{*}$ (with high probability). In the end, the first coordinate is refined to $\hat{\omega}_{1}=y(x_{1})/(\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1})$.
@@
where $w_{1}$ is already refined. Once more, using the probabilistic bisection process, the value of $z_{2}^{\top}\boldsymbol{\theta}^{*}$ is iteratively refined to $y(x_{2})$. Upon termination of the probabilistic bisection process, we obtain the aligned value of the second coordinate as
@@
Subsequently, at the $k^{th}$ iteration, we test on the sample $\boldsymbol{z}_{k}=\boldsymbol{\alpha}_{k}+\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$, where the true value equals
@@
After the value of $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}$ is refined to $y(x_{k})$ through Algorithm 2, we can obtain the aligned value of the $k^{th}$ coordinate as
@@
By asking a human to evaluate $x_{1},\cdots,x_{s}$ and by refining $\omega_{1},\cdots,\omega_{s}$ in sequence, we can refine the parameter $\boldsymbol{\theta}:=\sum_{i=1}^{s}\hat{\omega}_{i}\boldsymbol{\alpha}_{i}$ within a certain accuracy level. We include the pseudo code in Algorithm 4.
Note that the set of orthogonal bases is not unique; multiple sets may exist, and recognizing this possibility is essential. Consequently, we can conduct multiple rounds of alignment, with the option to switch the basis from one round to another.
@@
6:     Query sample $x_{k}$ through Algorithm 2 and get output $y(x_{k})$
@@ 7 Numerical Experiment
7.1 Synthetic Data Experiments
In the first experiment, we compare the performance of SL+LHF and SL using the same sample size to assess how parameter variations affect the performance of the two algorithms.
Experiment setup. Consider a high-dimensional linear regression problem where $d=100$ and $s=10$. The first ten coordinates of the true parameter are independently drawn from the uniform distribution, and the rest of the coordinates are zero. The observational noise is drawn from Gaussian distribution with a mean of zero and variance $\sigma\in\{1,2,5\}$; we choose $\kappa\in[0,0.95]$, where $\kappa$ comes from Assumption 4.4.1. Specifically, the difference of the utility in the two answers $c_{\Delta}^{+}(\theta)$ and $c_{\Delta}^{-}(\theta)$ is assumed to be $\|\theta-\theta^{*}\|^{\kappa}$. The selection precision (along each dimension) is set at $\varepsilon=0.1$, and the human expertise level in the choice model is $\gamma=1$. We conducted 30 repetitions for each experiment.
In the first stage of the two-stage framework, we selected a dataset with a sample size proportional to $\sigma^{2}\log d$ to input into Lasso. According to Theorem 5.1, this selection ensures that important features are identified with high probability. We chose the regularization parameter using cross-validation. The second stage is then executed, requiring human comparison samples, to reach the desired precision threshold $\varepsilon$. Once the SL+LHF algorithm completes, the combined sample size for both stages is recorded as the sample size for SL+LHF. Then, for a fair comparison, we provided the same sample size to SL (to train a lasso with the same number of observations) and computed the estimation error. We then looked at the ratio of the estimation errors of the two approaches, denoted by Err(SL+LHF)/Err(SL).
@@ 7.1 Synthetic Data Experiments
Figure 3 shows the median of the ratio Err(SL+LHF)/Err(SL) using different values of $\kappa$ when $\sigma$ takes values $1$, $2$, or $5$. For all values of $\sigma$, when $\kappa$ is smaller, the human’s selection accuracy is higher as $\theta$ approaches $\theta^{*}$. In particular, when $\kappa=0$, the selection accuracy is bounded away from 50%. Therefore, the task is harder as $\kappa$ increases. It is consistent with the increasing trend in the estimation error ratio as $\kappa$ grows from 0 to 1. In addition, a larger value of $\kappa$ leads to a larger sample size, which leads to a decrease in the estimation error of lasso. When $\sigma=1$, the intersection point of the curve, with the horizontal line of the ratio equal to $1$ (red dashed line), is around $\kappa=0.2$. That is, when $\kappa\leq 0.2$, SL+LHF outperforms pure SL in more than half of the cases. In particular, when $\kappa=0$, the median of the error ratio is close to 0.
Figure 3: Error ratio of the two-stage framework to the pure SL with different values of $\sigma$.
The results for noise levels with $\sigma=2$ and $\sigma=5$ are also illustrated in Figure 3. Notably, the trends in the error curves are similar. The threshold, defined as the intersection point with the red dashed line, is approximately 0.4 when $\sigma=2$ and approximately $0.65$ when $\sigma=5$.
Overall, the three curves show that the superiority of SL+LHF becomes more pronounced in the presence of higher observational noise. In conclusion, SL+LHF is more effective under increased noise or reduced $\kappa$.
7.1.2 The variation of expertise level $\gamma$.
We vary the parameter $\gamma$ to model different levels of human expertise and to analyze how the accuracy of human comparisons affects the estimation error. For the experiment setup, we set $\sigma=2$ and $s=10$. The smaller value of $\gamma$ indicates higher accuracy in the comparison task. In particular, as $\gamma$ approaches 0, the human can always select the accurate answer. Figure 4 shows the performance of the two algorithms with $\gamma$ equal to $0.8$, $1$, and $1.1$. Comparing the three curves, when $\gamma=1.1$, SL+LHF performs better than pure SL if $\kappa$ is lower than approximately 0.3; when $\gamma=0.8$, SL+LHF performs better than pure SL if $\kappa$ is lower than approximately 0.5. Therefore, we find support for our theory that the performance of SL+LHF improves when the human has a higher expertise level in the selection (i.e., the selection accuracy is higher).
Figure 4: Error ratio of the two-stage framework to the pure SL with different values of $\gamma$.
@@
The relative performance of two algorithms also depends on the dimension of the problem. For the experiment setup, we set $\gamma=1$ and $\sigma=2$. Figure 5 illustrates the algorithmic performance when important feature dimension $s$ takes values $10$, $20$, and $50$. The figure reveals that the intersection point shifts to a smaller value as the important feature dimension value increases. This observation aligns with our theoretical condition (5.2), indicating that as the value of the important feature dimension grows, the advantage of SL+LHF diminishes. This result highlights the importance of low-dimensional representations in the success of this method.
Figure 5: Error ratio of the two-stage framework to the pure SL with different values of $s$.
7.2 MTurk Experiment on the LNCA Ratio
In this section, we first outline the setup of our MTurk experiment [^3] https://www.mturk.com/. We then describe the estimation process for key parameters, including the accuracy convergence rate $\kappa$, expertise level $\gamma$, and noise level $\sigma$. After these parameters were estimated, we conducted simulations to compare the performance of the two-stage framework with that of the pure SL approach.
@@ 7.2 MTurk Experiment on the LNCA Ratio
We designed a task in which participants are asked to estimate the number of points within a square. We divided participants into two distinct groups. The first group was presented with one of the squares and two options for the estimated number of points. The second group was shown the same square and asked to provide their estimated number of points. (See Figures 6 and 7.) The task for the second group was to select the option they believed was closer to the actual number. For instance, in Figure 6(b), the true number is 32. Human participants were provided with two options: 55 and 35; the correct choice is 35.
@@
With this dual-group approach, we aimed to explore the efficacy of human estimation when participants were asked to estimate the number of points versus to choose from comparative options. In the latter case, we varied the discrepancy between the two choices to assess how the ease or difficulty of comparison affected the accuracy of participants’ selections. A larger disparity between the choices facilitated easier decision-making, while closer options posed a more challenging comparative task. Figure 6 shows three testing samples. Notably, sample (c) presents the most challenging comparison because of the minimal disparity ($|u(c^{+})-u(c^{-})|)$ observed between them. [^4] The gap between the first choice (45) and the true value (36) is 9 while the gap between the second choice (25) and the true value is 11. Thus, the difference between 9 and 11 is only 2.
To compare the performance of the two-stage framework and the pure supervised learning, we first estimated $\kappa$ and $\sigma$ from the online experiment in Amazon MTurk. Then we ran the simulation to compare two methods under different precision levels. We launched 50 different tasks, all structured in the same way as the task shown in Figure 7, and collected 500 data points in total. We estimated $\kappa$ and $\sigma$ as described in the following.
(a) 96 vs. 76. True: 64.
(b) 55 vs. 35. True: 32.
(c) 45 vs. 25. True: 36.
Figure 6: Experiment design for Group 1.
(a) Estimate the number of points.
Figure 7: Experiment design for Group 2.
@@
Estimation process for $\kappa$. Given figures with true parameter value $\theta^{*}$ (i.e., the number of points in the figure), we estimate the accuracy convergence rate $\kappa$ between two options using a human choice model. We collect multiple samples for each figure and use these estimates to derive the relationship between the difference in utility and the accuracy of human choices.
@@
Specifically, among two choices, assume the smaller one is $c^{-}$, the larger one is $c^{+}$, and the middle point is $\theta=(c^{-}+c^{+})/2$. According to the choice model, the human makes the correct choice with accuracy $p(\theta)=\frac{1}{1+\exp(-|u(c^{-})-u(c^{+})|/\gamma)}$. Note that we try to estimate $\kappa$ from Assumption 4.4.1 that
@@
That is, the accuracy is $p(\theta)=\frac{1}{1+\exp(-\lambda_{\Delta}/\gamma\cdot|\theta-\theta^{*}|^{\kappa})}$. Let $\tilde{\lambda}=\lambda_{\Delta}/\gamma$. For each data point with tested point $\theta_{i}$ and true parameter $\theta_{i}^{*}$, we define $y_{i}=1$ (the correct choice) with probability $p_{i}(\theta)$, and we define $y_{i}=0$ (the wrong choice) with probability $1-p_{i}(\theta)$. By maximizing the log-likelihood, we solve the optimization problem
$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i}),$
@@
$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}-y_{i}\log(1+\exp(-\tilde{\lambda}|\theta_{i}-\theta^{*}|^{\kappa}))-(1-y_{i})\log(1+\exp(\tilde{\lambda}|\theta_{i}-\theta_{i}^{*}|^{\kappa}))\,.$
We solve this optimization problem simply by enumerating different values of $\tilde{\lambda}$ and $\kappa$.
@@
Estimation process for $\sigma$. The other group is asked to directly estimate the percentage of total area. For the figure with true parameter values $\theta_{i}^{*}$, the human provides responses by $y_{i}=\theta_{i}^{*}+\epsilon_{i}$. We estimate the variance of the noise by $\hat{\sigma}^{2}=\sum_{i=1}^{N}\frac{(y_{i}-\theta_{i}^{*})^{2}}{N-1}$.
Empirical result. We applied a bootstrap resampling method to the data collected from Amazon MTurk, generating 500 resamples to estimate key parameters, including $\kappa$, $\tilde{\lambda}$, and $\sigma$. The resulting mean estimates are $\kappa=0.328$, $\tilde{\lambda}=0.132$, and $\sigma^{2}=29.665^{2}$. The variances are 0.061, 0.006, and 4.600 for $\kappa$, $\tilde{\lambda}$, and $\sigma$, respectively.
Figure 8 presents the median empirical error of the two-stage framework across varying precision levels, based on 100 repetitions for each value of $\varepsilon$. The results demonstrate that, for any given precision level, the empirical error remains within the targeted threshold. In addition, an upward trend in the empirical error is observed as the precision level increases. The green plot illustrates the sample size required to achieve a specified precision, showing a rapid decline in sample size as precision increases.
For a clearer comparison with the pure SL, Figure 9 displays the error ratio between the two-stage framework and the pure SL. When $\varepsilon=0.3\%$, the median of the estimation error of the two-stage framework is approximately 16% of the pure SL. Increasing the precision to $\varepsilon=4.3\%$, the median of the estimation error of the two-stage framework is approximately 45% of the pure SL. This trend of a decreasing error ratio with tighter precision aligns with Proposition 5.4, highlighting the increasing advantage of the two-stage framework as $\varepsilon$ approaches 0.
Figure 8: Empirical error of the two-stage framework for different values of $\varepsilon$.
Figure 9: Error ratio of the two-stage framework to the pure SL for different values of $\varepsilon$.
In conclusion, we acknowledge limitations in our numerical experiments. Ideally, our two-stage framework could be used to acquire responses sequentially from freelancers, first asking them which number is closer and and then asking them for an estimate of the number of points. Then the two results could be compared. However, implementing this comparison would be challenging because of the adaptive nature of our acquisition process, where each subsequent question depends on previously collected answers. Moreover, acquiring responses from the Amazon MTurk platform is costly. Given the complexity and expense of this procedure, we opted to use simulations to evaluate the performance of our framework.
8 Discussions and Concluding Remarks